OpenStudy (anonymous):

explain the difference between using the tangent ratio to solve for a missing angle in a right triangle versus using the cotangent ratio. Yo…

OpenStudy (anonymous):

tan theta = y/x cot theta = x/y Can you explain it from there?

OpenStudy (anonymous):

It's the reciprocal right? I kind of get it but the question says explain the difference of using one or the other with complete sentences and provide an example and it's hard to explain for me because I'm confused with trigonometric functions

OpenStudy (anonymous):

Like its hard to put into words

OpenStudy (anonymous):

Yes, they are reciprocals